# On the Generalization of Preference Learning with DPO

### v2

@@ 3.2 Reward Dynamics
We now analyze how the reward margin changes over the course of training, which allows us to bound the generalization error. We denote the input prompt as $x=(x^{(1)},x^{(2)},\dots,x^{(T)})$, where $x^{(i)}$ is the $i$-th token in the prompt and $T$ is the length of the prompt. We define the model output to be $f_{\theta}(x)=\text{softmax}(Wg(x))$, where $g:\mathcal{V}^{T}\mapsto\mathbb{R}^{d}$ is the mapping from the prompt to the final hidden state after normalization, and $W\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the unembedding layer matrix. The feature backbone can be either tunable or fixed, though the latter design choice is more relevant in practical applications. For example, in recently popularized parameter-efficient fine-tuning paradigm, the feature backbone is often kept frozen to prevent overfitting, and in black-box fine-tuning scenarios where the backbone is not exposed to the end-user.
In what follows, we first focus on a fixed encoder as a pragmatic approach to manage tractability while still extracting valuable insights into preference learning. This allows us to capture complex dynamics, which offers a clearer interpretation of the behaviors we aim to study. Later we will also investigate whether our theoretical insights hold when performing full fine-tuning, where the feature map is allowed to change. To follow the reward margins during training, we begin by deriving the dynamics of the weight matrix $W$ under gradient flow:
@@
where $W_{0}$ is the initial weight in the reference policy $\pi_{\text{ref}}$. $\tau$ determines the rate of change, where a larger $\tau$ corresponds to a slower rate of change. To ensure clarity in our exposition and elucidate the key insight, we first illustrate the derivation when the preferred response $y_{w,i}$ and non-preferred response $y_{l,i}$ consist of a token, encoded by the one-hot vector $\mathbf{y}_{w/l,i}$ in $\mathbb{R}^{|\mathcal{V}|}$. Our analysis will be expanded to a more complex multi-token setting in Section 4. Let $\Delta W=W-W_{0}$, a constant offset from $W$, we have:
@@
which contains the term of the reward margin. Since $\beta,\mathbf{y}_{w,j},\mathbf{y}_{l,j},x_{j}$ are fixed, we can consider the flow of the reward margin by multiplying $\beta(\mathbf{y}_{w,j}-\mathbf{y}_{l,j})^{\top}$ on the left and multiplying $g(x_{j})$ on the right of $\tau\Delta\dot{W}$. This yields of the dynamics for the reward margin:
@@ 4 Generalization Guarantees
4.1 Characterizing the Preference Distribution
We characterize the preference distribution by modeling the input feature to the unembedding layer. Importantly, the features we model are designed to reflect the characteristics of the real-world transformer backbone, ensuring that our theoretical analysis remains grounded in the specific inductive biases and structures that are typical of such models (see careful verification in Section 5). Specifically, we consider a preference distribution that consists of $K$ pairs of clusters that correspond to different concepts. In the context of alignment, the concepts can be broadly associated with different personality traits, political views, moral beliefs, etc. For example, the concepts may encompass common properties such as helpfulness, honesty, and harmlessness [21], and can also represent much more diversified and nuanced ones like conscientiousness, non-racism, compassion, and so on. For each concept, we have a pair of clusters containing samples aligned vs. misaligned with that concept.
@@ 4.2 Results
Given $Z\leq\min\left(\frac{1}{4l_{b}^{2}},Q^{1/4}-2\right)$, $d\leq 5Q$, $v\leq\frac{1}{4\sqrt{Q}}$, with probability at least $1-8KQ^{9/4}\exp\left(-\min\left(\frac{c\sqrt{Q}}{5},\frac{Q^{3/4}}{256}\right)\right)$ for some constant $c>0$, the trajectory $r_{i}(t)$ for all $i\in[N]$ is upper bounded by $r^{U}(t)$ and lower bounded by $r^{L}(t)$ which are given by
$r^{L}(t)=\frac{Q\beta^{2}}{4N\tau}t\quad r^{U}(t)=\frac{10Q\beta^{2}}{N\tau}t$
for $t\leq\tau_{1}=\frac{N\tau\log 3}{10Q\beta^{2}}$ and at $\tau_{1}$, for any training sample $\frac{\log 3}{40}\leq r(t)\leq\log 3$.
Interpretation.
@@
Building on our guarantee on the reward margin of training samples along with the fact that the reward dynamics of a new sample is of the same form as that of the training samples (cf. Equation (9)), we can bound the generalization error of the DPO reward model on the preference distribution.
@@
The generalization guarantee uses the fact that samples seen in training are predicted correctly to ensure that a new sample from the distribution is also likely to be classified correctly. This implies that the conditions for generalization are similar to those needed to guarantee strong training performance, which means less interaction between different types of preferences and a smaller number of clusters would also benefit the generalization error. In order to have less interaction between types of preferences or clusters, it would be necessary for the cluster directions to have smaller inner products which are only possible for a large number of clusters when the dimension is sufficiently large. This points to one reason as to why an increase in scale can allow for better model capabilities. Another aspect of the guarantees to consider is that they are for samples within the training distribution. As we see in Equation (9), the model behavior on new samples depends on the correlations between the new sample and its training samples, which may not be meaningful if the new sample is not well represented in the training set and can lead to unexpected behavior. This suggests that increasing scale and diversity of data can bolster a model’s ability to generalize.
@@ 4.3 Extension to Multi-Token Generation
To clearly see how the reward evolves and how each token contributes to the reward, we can decompose the reward for the $i$-th sample into the sum of token-wise rewards:
$r(y_{w/l,i})=\sum_{j=1}^{L}r(y_{w/l,i}^{(j)})=\sum_{j=1}^{L}\beta\log\frac{\pi_{\theta}(y_{w/l,i}|x_{i})}{\pi_{\text{ref}}(y_{w/l,i}|x_{i})},$ (12)
where $L$ is the length of the response, $y_{w/l,i}^{(j)}$ is the $j$-th token of a response to input $x_{i}$, and we use the subscript $w/l$ to indicate either preferred or non-preferred responses. Further, the likelihood of a response is given by $\pi_{\theta}(y_{w/l,i}|x_{i})=\prod_{j=1}^{L}p_{\theta}(y_{w/l,i}^{(j)}|x_{i},y_{w/l,i}^{(1)},...,y_{w/l,i}^{(j-1)})$, hence the token-wise reward can be expressed as:
@@ 5 Experimental Verification
We verify that our data assumption in Section 4 matches closely the characteristics of real-world alignment datasets. We consider the Anthropic Persona dataset [24], which well suits our study for two main reasons. First, the dataset is designed to capture a wide range of 135 behavioral styles and preferences, which allows us to validate our theorem under diverse preference distributions. Moreover, the persona dataset closely matches the theoretical setup, allowing us to define clusters concretely. Specifically, each persona has 500 statements that align and 500 statements that misalign with the persona trait, which can be viewed as a pair of concept clusters in our preference distribution. For instance, a persona “agreeableness” entails preferred statements like “It is important to treat other people with kindness and respect” that represents the persona, and also misaligned statements, e.g., “I tend to enjoy getting into confrontations and arguments with others”.
Recall that the data distribution under which our results hold is that (1) the embeddings consist of a shared component along some direction and (2) each concept or cluster varies along orthogonal directions. To verify the shared component, we compute the average cosine similarity between the final embeddings of statements from different pairs of personas. The embeddings are extracted from the LlaMa-2-7B model [6], a popular open-source foundation model with accessible internal representations. As depicted in Figure 1(a), the average similarity is high, confirming the shared structure among a random subset of 10 personas. Furthermore, to verify the orthogonality assumption, we subtract the shared component from each embedding vector, and then compute the average cosine similarity for any pair of personas. As seen in Figure 1(b), the average cosine similarity is close to 0 for non-diagonal entries, suggesting the remaining components are nearly orthogonal. For completeness, we provide verification across all personas in Appendix C.
Figure 2: Visualization of average reward margins over the course of training across a different number of clusters.
@@
In Theorem 4.1, we show that the rate at which the reward margin increases, $\dot{r}$, decreases as the number of clusters or concepts increases in training. To verify this empirically, we randomly sample different numbers of personas from the Anthropic dataset, simulating the varying number of concepts $K=\{1,2,4,8,16\}$. For each setting, we perform full fine-tuning on the LLaMA-2 model [6] using the DPO loss. As depicted in Figure 2(a), the training reward margin grows more rapidly for smaller $K$, given the same number of training steps. Similarly, we verify our Theorem 4.2 in Figure 2(b), which shows that the test reward margin on new inputs also exhibits the same trend. Moreover, we find a similar decrease in the rate at which the loss and accuracy change and provide results in Appendix C. These results validate that our theoretical insights indeed translate to practical alignment process.
6 Future Extension Beyond DPO
@@ 7 Related Works
A key aspect of training and deploying large language models is ensuring the models behave in safe and helpful ways [27, 28, 29, 30]. This is an important problem due to the potential harms that can arise in large models [31, 32, 24, 33, 34, 35, 36, 37, 38, 39, 40]. A wide range of methods have been developed that utilize human feedback or human preference data to train models to avoid harmful responses and elicit safer or more helpful responses [19, 22, 41, 42, 20, 21, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Particularly, the Reinforcement Learning from Human Feedback (RLHF) framework has proven effective in aligning large pre-trained language models [19, 22, 20, 21]. However, given its computational inefficiency, recent shifts in focus favor closed-form losses that directly utilize offline preferences, like Direct Preference Optimization (DPO) [18] and related methodologies [55, 56, 57, 58, 59, 25, 26, 60, 61, 62, 63, 64, 65, 66, 67, 68]. Despite the empirical success and wide adoption in real-world systems [4, 5, 6], fewer works provide theoretical underpinnings [55, 69, 70, 25, 64, 71, 72, 73, 74]. In this work, we make an initial attempt to comprehensively analyze the generalization behavior of preference optimization from a rigorous theoretical standpoint. We introduce a new theoretical framework specifically designed to examine the generalization properties of LLMs by approximating their reward dynamics, providing insights into practical aspects of aligning LLMs.
@@
Understanding how and why deep models generalize has been a subject of extensive research. One approach is through the lens of feature learning, attempting to understand how models learn data-dependent features and how these features are structured [75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. Another approach is through providing generalization bounds that quantify the expected performance of the model beyond the training samples and over a data distribution [11, 12, 13, 14, 15, 16, 17, 7, 8, 10, 9]. While existing generalization theories typically consider simpler learning tasks such as regression and classification, our work provides generalization analysis in the context of aligning language models, which entails dealing with the complex output space of sentences. Moreover, existing generalization theory typically considers overparameterized models that achieve near-optimal loss [11, 12, 13, 14] or are independent of the training process [15, 16, 17]. In contrast, our framework focuses on the generalization of models by analyzing the reward dynamics after finite gradient steps, which matches more closely with the real-world practices of aligning LLMs. Our theoretical insights are further supported by empirical validations on contemporary LLMs, as shown in Section 5.
@@ 8 Conclusion
Our work theoretically analyzes the generalization behavior of preference learning, which remains an open problem in the field of AI safety. We base our theoretical analysis on a popular alignment loss, direct preference optimization, which implicitly learns a reward model. Key to our framework, we analyze the reward margin associated with each sample and its trajectory throughout the training process, which allows us to effectively bound the generalization error. Through rigorous analysis, we establish conditions under which the model trained with DPO loss generalizes to new inputs with provably high accuracy. Empirical validation on contemporary LLMs and real-world alignment datasets confirms the practical relevance of our framework, offering insights crucial for developing AI systems that align with human intentions and preferences. We hope our work catalyzes future investigations into the theoretical understanding of alignment methods.
@@ Acknowledgement
$\mathcal{L}_{0-1}(\mathcal{P})\leq 2KQ^{2}e^{-\epsilon^{2}/2(2+dv^{2}+\epsilon v)}$ (45)
@@
$\mathcal{L}_{0-1}(\mathcal{P})\leq 2KQ^{2}e^{-\epsilon^{2}/2(2+dv^{2}+\epsilon v)}$ (57)
@@ Appendix B Multi-Token Derivation
Appendix C Additional Verification
@@
Here we provide the plot of the cosine similarities of embeddings between different personas before and after subtracting the mean embedding in Figure 3(a) and 3(b). The personas are ordered according to lexicographical order.
@@
We present the training and test losses and accuracies across different numbers of clusters as seen in Figures 4(a), 4(b), 5(a), and 5(b). We find that the losses decrease at a slower rate and the accuracies increase at a slower rate as the number of clusters increase.
@@
Appendix D Training and Experimental Details

### v3

@@ 4.2 Reward Dynamics
Our theory revolves around analyzing how the reward margin changes over the course of training, which allows us to bound the generalization error after finite-step DPO updates.
Dynamics of weight.
We begin by analyzing the dynamics of the weight matrix given the latent representation of samples. A standard setup for training is to apply gradient descent, in which case, the dynamics of the weight matrix $W$ at step $t$ is:
$W(t+1)-W(t)=\frac{\eta}{N}\sum_{i=1}^{N}\beta\sigma(-\beta(\mathbf{y}_{w,i}-\mathbf{y}_{l,i})^{\top}(W(t)-W_{0})g(x_{i}))(\mathbf{y}_{w,i}-\mathbf{y}_{l,i})g(x_{i})^{\top},$ (5)
where $W_{0}$ is the initial weight in the reference policy $\pi_{\text{ref}}$ and $\eta$ is the learning rate. We consider for our theoretical analysis, gradient flow, a continuous approximation of gradient descent. To follow the reward margins during training, we derive the dynamics of the weight matrix $W$ under gradient flow:
@@
where $\tau$ determines the rate of change, where a larger $\tau$ corresponds to a slower rate of change. To ensure clarity in our exposition and elucidate the key insight, we first illustrate the derivation when the preferred response $y_{w,i}$ and non-preferred response $y_{l,i}$ consist of a token, encoded by the one-hot vector $\mathbf{y}_{w/l,i}$ in $\mathbb{R}^{|\mathcal{V}|}$. Our analysis will be expanded to a more complex multi-token setting in Section 5.
Dynamics of reward margin.
Let $\Delta W=W-W_{0}$, a constant offset from $W$, we have:
@@
which contains the term of the reward margin. Since $\beta,\mathbf{y}_{w,j},\mathbf{y}_{l,j},x_{j}$ are fixed, we can consider the flow of the reward margin by multiplying $\beta(\mathbf{y}_{w,j}-\mathbf{y}_{l,j})^{\top}$ on the left and multiplying $g(x_{j})$ on the right of $\tau\Delta\dot{W}$. This yields the dynamics for the reward margin:
@@ 5 Generalization Guarantees
5.1 Characterizing the Diverse Preference Distribution
We characterize the preference distribution by modeling the input feature to the unembedding layer. Importantly, the features we model are designed to reflect the characteristics of the real-world transformer backbone, ensuring that our theoretical analysis remains grounded in the specific inductive biases and structures that are typical of such models (see empirical verification in Section 6). Specifically, we consider a preference distribution that consists of $K$ pairs of clusters that correspond to different concepts. In the context of alignment, the concepts can be broadly associated with different personality traits, political views, moral beliefs, etc. For example, the concepts may encompass common properties such as helpfulness, honesty, and harmlessness [21], and can also represent much more diversified and nuanced ones like conscientiousness, non-racism, compassion, and so on [24]. For each concept, we have a pair of clusters containing samples aligned vs. misaligned with that concept.
@@ 5.2 Results
Given $Z\leq\min\left(\frac{1}{4l_{b}^{2}},Q^{1/4}-2\right)$, $d\leq 5Q$, $v\leq\frac{1}{4\sqrt{Q}}$, with probability at least $1-8KQ^{9/4}\exp\left(-\min\left(\frac{c\sqrt{Q}}{5},\frac{Q^{3/4}}{256}\right)\right)$ for some constant $c>0$, the trajectory $r_{i}(t)$ for all $i\in[N]$ is upper bounded by $r^{U}(t)$ and lower bounded by $r^{L}(t)$ which are given by $r^{L}(t)=\frac{Q\beta^{2}}{4N\tau}t$ and $r^{U}(t)=\frac{10Q\beta^{2}}{N\tau}t$ for $t\leq\tau_{1}=\frac{N\tau\log 3}{10Q\beta^{2}}$ and at $\tau_{1}$, for any training sample $\frac{\log 3}{40}\leq r(t)\leq\log 3$.
Theoretical insights.
@@
Building on our guarantee on the reward margin of training samples along with the fact that the reward dynamics of a new sample is of the same form as that of the training samples (cf. Equation (9)), we can bound the generalization error of the DPO reward model on the preference distribution. We present a simplified bound for clarity, and provide a tighter bound in Appendix A.
@@
The generalization guarantee uses the fact that samples seen in training are predicted correctly to ensure that a new sample from the distribution is also likely to be classified correctly. This implies that the conditions for generalization are similar to those needed to guarantee strong training performance, which means less interaction between different types of preferences and a smaller number of clusters would also benefit the generalization error. In order to have less interaction between types of preferences or clusters, it would be necessary for the cluster directions to have smaller inner products which are only possible for a large number of clusters when the dimension is sufficiently large. This points to one reason as to why an increase in scale can allow for better model capabilities. Another aspect of the guarantees to consider is that they are for samples within the training distribution. As we see in Equation (9), the model behavior on new samples depends on the correlations between the new sample and its training samples, which may not be meaningful if the new sample is not well represented in the training set. This suggests that increasing scale and diversity of data can bolster a model’s ability to generalize.
@@ 5.3 Extension to Multi-Token Generation
To clearly see how the reward evolves and how each token contributes to the reward, we can decompose the reward for the $i$-th sample into the sum of token-wise rewards: $r(y_{w/l,i})=\sum_{j=1}^{L}r(y_{w/l,i}^{(j)})=\sum_{j=1}^{L}\beta\log\frac{\pi_{\theta}(y_{w/l,i}|x_{i})}{\pi_{\text{ref}}(y_{w/l,i}|x_{i})}$, where $L$ is the length of the response, $y_{w/l,i}^{(j)}$ is the $j$-th token of a response to input $x_{i}$, and we use the subscript $w/l$ to indicate either preferred or non-preferred responses. Further, the likelihood of a response is given by $\pi_{\theta}(y_{w/l,i}|x_{i})=\prod_{j=1}^{L}p_{\theta}(y_{w/l,i}^{(j)}|x_{i},y_{w/l,i}^{(1)},...,y_{w/l,i}^{(j-1)})$, hence the token-wise reward can be expressed as:
@@ 6 Empirical Verification
We verify that our data assumption in Section 5 matches closely the characteristics of real-world alignment datasets. We consider the Anthropic Persona dataset [24].Recall that the data distribution under which our results hold is that (1) the embeddings consist of a shared component along some direction and (2) each concept or cluster varies along orthogonal directions. To verify the shared component, we compute the average cosine similarity between the final embeddings of statements from different pairs of personas. The embeddings are extracted from the LlaMa-2-7B model [6], a popular open-source foundation model with accessible internal representations. As depicted in Figure 3(a), the average similarity is high, confirming the shared structure among a random subset of 10 personas. Furthermore, to verify the orthogonality assumption, we subtract the shared component from each embedding vector, and then compute the average cosine similarity for any pair of personas. As seen in Figure 3(b), the average cosine similarity is close to 0 for non-diagonal entries, suggesting the remaining components are nearly orthogonal. For completeness, we provide verification across all personas in Appendix D.
Figure 4: Average reward margins over the course of training across a different number of clusters.
@@
In Theorem 5.1, we show that the rate at which the reward margin increases, $\dot{r}$, decreases as the number of clusters or concepts increases in training. To verify this empirically, we randomly sample different numbers of personas from the Anthropic dataset, simulating the varying number of concepts $K=\{1,2,4,8,16\}$. For each setting, we perform full fine-tuning on the LLaMA-2 model [6] using the DPO loss. As depicted in Figure 4(a), the training reward margin grows more rapidly for smaller $K$, given the same number of training steps. Similarly, we verify our Theorem 5.2 in Figure 4(b), which shows that the test reward margin on new inputs also exhibits the same trend. Moreover, we find a similar decrease in the rate at which the loss and accuracy change and provide results in Appendix D. These results validate that our theoretical insights indeed translate to practical alignment process.
@@ 7 Related Works
A key aspect of training and deploying large language models is ensuring the models behave in safe and helpful ways [25, 26, 27, 28]. This is an important problem due to the potential harms that can arise in large models [29, 30, 24, 31, 32, 33, 34, 35, 36, 37, 38]. A wide range of methods have been developed that utilize human feedback or human preference data to train models to avoid harmful responses and elicit safer or more helpful responses [19, 22, 39, 40, 20, 21, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Particularly, the Reinforcement Learning from Human Feedback (RLHF) framework has proven effective in aligning large pre-trained language models [19, 22, 20, 21]. However, given its computational inefficiency, recent shifts in focus favor closed-form losses that directly utilize offline preferences, like Direct Preference Optimization (DPO) [18] and related methodologies [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. Despite the empirical success and wide adoption in real-world systems [4, 5, 6], fewer works provide theoretical underpinnings [53, 69, 70, 58, 64, 71, 72, 73, 74]. In this work, we make an initial attempt to comprehensively analyze the generalization behavior of preference optimization from a rigorous theoretical standpoint. Our work considers offline preference optimization which differs from the setting of other theoretical works on preference-bases reinforcement learning [75, 76]. We introduce a new theoretical framework specifically designed to examine the generalization properties of LLMs by approximating their reward dynamics, providing insights into practical aspects of aligning LLMs.
@@
Understanding how and why deep models generalize has been a subject of extensive research. One approach is through the lens of feature learning, attempting to understand how models learn data-dependent features and how these features are structured [77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. Another approach is through providing generalization bounds that quantify the expected performance of the model beyond the training samples and over a data distribution [11, 12, 13, 14, 15, 16, 17, 7, 8, 10, 9]. While existing generalization theories typically consider simpler learning tasks such as regression and classification, our work provides generalization analysis in the context of aligning language models, which entails dealing with the complex output space of sentences. Moreover, existing generalization theory typically considers overparameterized models that achieve near-optimal loss [11, 12, 13, 14] or are independent of the training process [15, 16, 17]. One line of works considers algorithmic stability which allows for generalization bounds that are dependent on the number of steps [87, 88]. In contrast, our framework focuses on the generalization of models by directly following and analyzing the reward dynamics after finite gradient steps, which matches more closely with the real-world practices of aligning LLMs. Our theoretical insights are further supported by empirical validations on contemporary LLMs, as shown in Section 6.
@@ 8 Conclusion
Our work theoretically analyzes the generalization behavior of preference learning, which remains an open problem in the field of AI safety. We base our theoretical analysis on a popular alignment loss, direct preference optimization, which implicitly learns a reward model. Key to our framework, we analyze the reward margin associated with each sample and its trajectory throughout the training process, which allows us to effectively bound the generalization error. Through rigorous analysis, we establish conditions under which the model trained with DPO loss generalizes to new inputs with provably high accuracy. Empirical validation on contemporary LLMs and real-world alignment datasets confirms the practical relevance of our framework, offering insights crucial for developing AI systems that align with human intentions and preferences. We hope our work catalyzes future investigations into the theoretical understanding of preference optimization methods.
9 Limitations
While our study primarily focuses on DPO as a representative case, it is important to acknowledge that our analysis may not fully capture the nuances of other emerging preference learning methods. We envision that our theoretical framework and insights can be extended to these methods, which we discuss in Appendix C. Future work should investigate the applicability and adaptability of our framework to these newer approaches, ensuring a comprehensive understanding of generalization across a broader spectrum of preference learning methodologies.
@@ Acknowledgement
$\mathcal{R}(\mathcal{P})\leq 2KQ^{2}e^{-\epsilon^{2}/2(2+dv^{2}+\epsilon v)}$ (42)
@@
$\mathcal{R}(\mathcal{P})\leq 2KQ^{2}e^{-\epsilon^{2}/2(2+dv^{2}+\epsilon v)}$ (54)
@@ Appendix B Multi-Token Derivation
Appendix C Future Extension Beyond DPO
Appendix D Additional Verification
@@
Here we provide the plot of the cosine similarities of embeddings between different personas before and after subtracting the mean embedding in Figure 5(a) and 5(b). The personas are ordered according to lexicographical order.
@@
Gaussian Cluster Verification
We verify that the cluster component of embeddings from real-world models and datasets can reasonably be modeled by a Gaussian distribution. We use the Anthropic Persona dataset [24] which consists of a diverse set of personas. For each persona, we collect the final layer embeddings at the end of each positive statement and normalize them to have unit norm on average. We calculate the average over personas of the Frobenius norm of the covariance matrix and the average squared distance from the mean of these embeddings. These are 0.058 and 0.227 respectively, suggesting that the overall variance is relatively small and a Gaussian distribution would be sufficient to capture the variance of the embedding distributions.
@@
We present the training and test losses and accuracies across different numbers of clusters as seen in Figures 6(a), 6(b), 7(a), and 7(b). We find that the losses decrease at a slower rate and the accuracies increase at a slower rate as the number of clusters increase.
@@
Verification on Llama-3.1-8B
We provide verification of the generalization results with the same training setup as with LlaMa-2-7B and provide the results in Figures 8(a), 8(b), 9(a), 9(b), 10(a), 10(b).
Figure 8: Llama-3.1-8B: Average reward margins over the course of training across a different number of clusters.
Figure 9: Llama-3.1-8B: Visualization of loss over the course of training across a different number of clusters.
Figure 10: Llama-3.1-8B: Visualization of accuracy over the course of training across a different number of clusters.
Appendix E Training and Experimental Details